# The Battle of the Water Futures

### v1

@@ 2.1.3 Water Sources
Water sources represent the entire process of water abstraction and treatment consolidated at a single location. Each source has a nominal production capacity which represents a hard limit on the amount of water that can be delivered in any 24-hour window.
In the BWF, three macro-types of water sources are modelled: groundwater, surface water, and desalination. The main trade-off is that groundwater sources are generally smaller (< 30 $Mm^{3}/year$) and cheaper to operate, as the water has higher quality and requires less treatment. Surface water treatment plants and desalination plants are generally much larger (30-60 $Mm^{3}/year$) but are considerably more expensive and energy-intensive to run. Moreover, surface water sources differ as they are affected by climate conditions; specifically, low inflows within rivers could temporarily shut down treatment plants downstream (effectively reducing their capacity to zero for those days). Instead, groundwater sources have an extraction permit (expressed in $m^{3}$ per year). This is not a hard physical constraint such as the nominal production capacity, but rather a “soft” legislative constraint checked by the government at the end of each year. This penalty effectively represents compensation for the hydrological displacement affecting farmers and natural areas due to the overextraction. The fine amount is set by law and can therefore change at any time based on political decisions.
The cost of water production at every source is a combination of four components:
  • Fixed costs (personnel, taxes, planned maintenance, etc.)
  • Volumetric costs for energy
  • Volumetric costs for non-energy related expenditures (chemicals, filters, etc.)
  • Extra volumetric costs for non-energy related expenditures This final component is an additional cost paid when the source operates above its planned production capacity, i.e., nominal capacity multiplied by its capacity target factor (which can be interpreted as the ideal efficiency point).
@@
New water sources have an uncertain construction time, so participants must communicate the construction start date, and the activation date will be randomized. The capital investment needed for a new source is simply the unit cost of construction at that year by the required size.
@@
Unit cost of construction Dynamic Endogenous Source Size Class $\times$ National $\texteuro/m^{3}\dot{d}ay$
Operational costs - fixed Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\texteuro/year$
Operational costs - volumetric for energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $kWh/m^{3}$
Operational costs - volumetric for non-energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\texteuro/m^{3}$
@@
Fine amount Dynamic Exogenous Groundwater source permit excedance Severity Class $m^{3}/year$
@@ 2.1.4 Pumping stations
Whenever a new source is opened, participants must define which pump option (selected exclusively from a pre-defined list of available pumps) and how many units are installed at that source’s pumping station. Similarly, participants can decide to replace pumps at existing locations.
@@
Property Type Scope Identifier Static Pumping station Assigned source Static Pumping station Pump option Static [Decision] Pumping station Number of pumps Static [Decision] Pumping station
@@
Property Type Scope
@@
Expected lifetime Static [Uncertain] Pump option
@@ 2.1.5 Connections (Pipes)
Connections are either completely within a province (intra-province) or shared between provinces (inter-province). The complete lists of intra-province and inter-province connections are included in the data files mapped in Appendix A.
@@
Property Type Scope Unit Identifier Static Connection Node A Static Connection Node B Static Connection Type Static Connection Distance Static Connection m (Minor loss coefficient) set to 0
@@
Darcy friction factor - decay rate Static [Uncertain] Pipe option
@@
Cost (new pipe) Dynamic endogenous Pipe option €/m
Equivalent emissions (new pipe) Dynamic exogenous Pipe option tCO2eq/m
@@ 2.2 System External Drivers
This chapter describes the external drivers that influence the water distribution system and how they have been modelled in the BWF competition. The key external systems modelled are: - the climate, - the energy system, and - the economic context.
@@ 2.2.2 Energy System
We will model the following three different parameters for the evaluation of the water distribution networks.
Variable Variable Type Presented By
Elecricity price in €/kWh Dynamic Exogenous National (Simulation model)
Greenhouse gas emission by electricity production in kgCO2/kWh Dynamic Exogenous National (Simulation model)
Unit costs for solar pv panel in €/kW Dynamic Exogenous National (Simulation model)
We assume yearly changes in electricity price roughly influenced by inflation but with a high uncertainty. Due to different electricity demands during the day, the electricity price will change hourly following a daily pattern with peaks during the working hours and lows during the night. For each kWh obtained from the power suppliers we will estimate the produced greenhouse gas emissions from generating the energy. This value may change yearly due to different compositions of the electricity generation. The alternative to buying electricity from a power supplier is to generate it from solar pv panels that also come with specific costs. These costs may change yearly due to more efficiency and different climate conditions.
The action the participants can do is to install solar pv panels on pump stations at any time, that will have a lifespan of 25 years. The main objective is to achieve low electricity costs in the water distribution network. As a secondary objective the greenhouse gas emissions produced by the water distribution network should be minimized.
@@ 2.2.3.1 Bond Issuance
The bond issuance governs how additional capital is raised when a utility exhausts its allocated budget for a period. When the available budget is fully utilized, a bond is automatically issued to cover the costs of planned interventions. The price and yield of this bond are determined by the modeled bond demand and market conditions specific to the issuance year. These bonds introduce future repayment obligations, and all associated interest payments accumulate as part of the total financial cost of the interventions. While utilities cannot directly control bond yields, they can anticipate debt accumulation through scenario analysis and adopt strategies that maintain financial sustainability. For instance, utilities may design pacing of interventions to minimize borrowing, or evaluate alternative interventions that reduce the likelihood of large bond issuances during periods with unfavorable yield conditions.
Variable Variable Type Presented By
Remaining budget in year t Dynamic Endogenous Utility ID
Bond issuance trigger (budget exhausted) Dynamic Endogenous Utility ID
Bond amount issued Dynamic Endogenous Utility ID
Bond price (issue price) Dynamic Exogenous Simulation model
Bond yield Dynamic Exogenous Simulation model
Bond demand curve Dynamic Exogenous Simulation model
Interest payments Dynamic Endogenous Utility ID
Total financial cost (capex + interest) Dynamic Endogenous Utility ID
2.2.3.2 Budget Allocation
The budget allocation module governs how national funding is distributed across provincial water utilities, each of which operates and maintains its own water distribution system. This allocation affects the resources available to implement interventions and support operations. Budget distribution can follow several principled rules: proportional allocation to population size; inverse-proportional schemes favoring smaller regions; income-based allocations reflecting socioeconomic considerations; inverse-proportional schemes favoring less wealthy regions; or fully customized schemes.
Variable Variable Type Presented By
National total budget Static National
Population size Dynamic Endogenous Municipality ID
Income index (socio-economic) Dynamic Endogenous Municipality ID
Per-capita allocation weight Option Rule
Inverse-population allocation weight Option Rule
Income-based allocation weight Option Rule
Equity priority score (favoring less wealthy regions) Option Rule
Customized allocation vector Option Participant-defined
2.2.3.3 Inflation Dynamics
Inflation represents the annual rate at which general price levels evolve over the planning horizon. In the BWF setting, inflation affects all monetary flows: capital expenditures, operating expenses, various costs, and prices for energy consumption and water tariffs. Because inflation compounds over time, it can significantly alter the real purchasing power of fixed budgets and shape the financial feasibility of long-term interventions. Interventions on inflation are limited, as inflation is an exogenous macroeconomic indicator outside the direct control of water utilities. However, utilities may adopt strategies that account for inflation, such as front-loading major investments to avoid future cost escalation, or stress-testing the interventions under alternative inflation trajectories. The primary goal is to design actions and financial flows that remain robust despite uncertain future price conditions.
Variable Variable Type Presented By
Inflation rate in year t Dynamic Exogenous National (Simulation model)
2.2.3.4 Water Tariffs
The water tariffs module models how utilities generate revenue through a pricing scheme commonly used in real-world systems: a fixed component (e.g., service charge) combined with a volumetric component (per-m³ charge). Both components evolve annually according to the prevailing inflation rate, ensuring that real revenue tracks changes in general price levels. Both the fixed and volumetric tariff components are exogenously given and evolve solely according to the uncertain inflation dynamics defined in the scenario generator. Participants must therefore design plans that remain financially robust under tariff paths they cannot control. This constraint emphasizes the importance of planning under deep uncertainty: utilities must cope with potential revenue volatility driven by inflation.
Variable Variable Type Presented By
Fixed tariff component (service charge) Dynamic Exogenous Utility ID
Volumetric tariff component [€/m³] Dynamic Exogenous Utility ID
Annual tariff escalation rule Dynamic Exogenous National (inflation-indexed)
2.3 System Interventions

### v4

@@ 2.1.3 Water Sources
Water sources represent the entire process of water abstraction and treatment consolidated at a single location. Each source has a nominal production capacity which represents a hard limit on the amount of water that can be delivered throughout any consecutive 24-hour window.
In the BWF, three macro-types of water sources are modelled: groundwater, surface water, and desalination. The main trade-off is that groundwater sources are generally smaller (< 25 $Mm^{3}/year$) and cheaper to operate, as the water has higher quality and requires less treatment. Surface water treatment plants and desalination plants are generally much larger (30-45 $Mm^{3}/year$) but are considerably more expensive and energy-intensive to run. Moreover, surface water sources differ as they are affected by climate conditions; specifically, low inflows within rivers could temporarily shut down treatment plants downstream (effectively reducing their capacity to zero for those days). Instead, groundwater sources have an extraction permit (expressed in $m^{3}$ per year). This is not a hard physical constraint such as the nominal production capacity, but rather a “soft” legislative constraint checked by the government at the end of each year. This penalty effectively represents compensation for the hydrological displacement affecting farmers and natural areas due to the overextraction. The fine amount is set by law and can therefore change at any time based on political decisions.
The cost of water production at every source $s$ is a combination of four components:
[…]
where:
  • $F_{s}$ represents fixed costs, including personnel, taxes, and planned maintenance.
  • $c^{\text{el}}(y)\cdot E^{\text{grid}}_{s}(y)$ represents volumetric costs for energy.
  • $c^{\text{vol}}_{s}(y)\cdot Q_{s}(y)$ represents the volumetric costs for non-energy related expenditures, such as chemicals and filters.
  • $c^{\text{extra}}_{s}(y)\cdot\max\bigl(Q_{s}(y)-\phi_{s}\cdot Q^{*}_{s},0\bigr)$ represent the extra volumetric costs incurred when production exceeds the planned threshold.
In Equation 5, $Q_{s}(y)$ is the total volume produced by the source in year $y$, $Q^{*}_{s}$ is the source nominal capacity, and $\phi_{s}$ is the capacity target factor (source-type dependent), which defines the ideal operating point above which additional costs are applied.
Additionally, the variable $E^{\text{grid}}_{s}(y)$ represents the electrical grid energy consumed for water production by the source $s$ in year $y$ and it is calculated as:
[…]
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $E_{s}(t)$ represents the energy consumed by the sources $s$ at time $t$, and the $E^{\text{PV}}_{s}(t)$ is electricity generated by the photovoltaic panels associated with the source $s$ at time $t$.
The source energy consumption $E_{s}(t)$ is:
[…]
where $\epsilon_{s}$ is the source specific energy and $Q_{s}(t)$ is the volume produced by the source at timestep $t$.
To complete the accounting, the total operational expenditure associated with all the sources in water utility $w$ at year $y$ is:
[…]
where $\mathcal{S}_{w}$ represents the collection of water sources managed by utility $w$, and $\mathrm{OPEX}_{s}(y)$ is the individual source cost defined in Equation 5.
@@
New water sources have an uncertain construction time, so participants must communicate the construction start date, and the activation date will be randomized.
The capital investment associated with the construction of new sources in water utility $w$ at year $y$ is:
[…]
where for a source $s$ in the set of the water utility’s sources ($\mathcal{S}_{w}$), $\tau_{s}$ is its starting construction time, $\mathbf{1}_{\{\tau_{s}=y\}}$ is an indicator function equal to 1 if the construction begins in year $y$ (0 otherwise), $Q^{*}_{s}$ is the requested source nominal capacity, and $c^{\text{source}}(\text{class}(s),y)$ the unit cost, which depends on the year of construction and the source class (see Table LABEL:tbl:sources-classes).
Table 6: Sources classification by their nominal capacity.
Source nominal capacity $Mm^{3}/year$] Source Class 4 SMALL 8 MEDIUM 16 LARGE > 16 VERY LARGE
@@
Unit cost of construction Dynamic Endogenous Source Size Class $\times$ National $\text{\texteuro}/(m^{3}/day)$
Operational costs - fixed Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\text{\texteuro}/year$
Specific energy Static [Uncertain] Source $kWh/m^{3}$
Operational costs - volumetric for non-energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\text{\texteuro}/m^{3}$
@@
Fine amount Dynamic Exogenous Groundwater source permit excedance Severity Class
@@ 2.1.4 Pumping stations
Whenever a new source is opened, participants must define which pump option (selected exclusively from a pre-defined list of available pumps) and how many units are installed at that source’s pumping station. Similarly, participants can decide to replace pumps at existing locations or add new ones of the same type as installed.
Whether replacing old pumps or being added in new pumping stations, the capital investment associated with the installation of pumps in water utility $w$ at year $y$ is:
[…]
where $n$ is the pumping station index within the set of the water utility’s pumping stations ($\mathcal{N}_{w}$), $p$ is the p-th pump in that pumping station set ($\mathcal{P}_{n}$), $\tau_{p}$ is the pump installation time, $\mathbf{1}_{\{\tau_{p}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), and $c^{\text{pump}}(\gamma_{p},y)$ the unit cost for the selected pump option $\gamma_{p}$.
@@
Thus, the operational expenditure associated with pumping in water utility $w$ at year $y$ is:
[…]
where $n$ is the pumping station index within the set of the water utility’s pumping stations ($\mathcal{N}_{w}$), $c^{\text{el}}(y)$ is the dynamic electricity price, and $E^{\text{grid}}_{n}(y)$ is the electricity consumption from the grid for the n-th pumping station in year $y$. This quantity is computed as:
[…]
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $E_{p}(t)$ the p-th pump energy consumption at time $t$, $\mathcal{P}_{n}$ the set of pumps in pumping station $n$, and $E^{\text{PV}}_{n}(t)$ is electricity generated by the photovoltaic panels associated with pumping station $n$ at time $t$.
@@
Property Type Scope Unit Identifier Static Pumping station Assigned source Static Pumping station Pump option Decision Pumping station Number of pumps Decision Pumping station unit
@@
Property Type Scope Unit
@@
Lifetime Static [Uncertain] Pump option years
Unit cost Dynamic Endogenous Pump option
@@ 2.1.5 Connections (Pipes)
Whether replacing old pipes or being added in a connection, the capital investment associated with the installation of pipes in water utility $w$ at year $y$ is:
[…]
where $\mathcal{J}_{w}$ is the set of all pipes within water utility $w$, $j$ the pipe index, $\tau_{j}$ is the pipe installation time, $\mathbf{1}_{\{\tau_{j}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), $c^{\text{pipe}}(\theta_{j},y)$ the unit cost for the selected pipe option $\theta_{j}$, and $L_{j}$ is the length of pipe $j$.
@@
Connections are either completely within a province (intra-province) or shared between provinces (inter-province). The complete lists of intra-province and inter-province connections are included in the data files mapped in Appendix A. The capital cost for inter-provincial pipes is shared equally between the water utilities operating in the connected provinces (i.e., Equation 13 is used but the pipe’s length is half).
@@
Property Type Scope Unit Identifier Static Connection Node A Static Connection Node B Static Connection Type Static Connection Distance Static Connection m Pipe option Decision Connection Pipe installation date Decision Connection (Minor loss coefficient) set to 0
@@
Darcy friction factor - decay rate Static [Uncertain] Pipe option $years^{-}1$
@@
Lifetime Static [Uncertain] Pipe option yearsyears
Unit cost (new pipe) Dynamic endogenous Pipe option $\text{\texteuro}/m$
Equivalent emissions (new pipe) Dynamic exogenous Pipe option $\text{tCO2eq}/m$
@@ 2.2 System External Drivers
This chapter describes the external drivers that influence the national WDS and how they have been modelled in the BWF competition. The key external systems modelled are the climate, the energy system, and the economic context.
@@ 2.2.1 Climate & Hydro-meteorological Forcing
Climate is an exogenous factor beyond the control of the national water sector. Four meteorological variables are relevant for the national grid planning: temperature, precipitation, solar radiation and evotranspiration.
The average maximum temperature affects water demand seasonality.
Temperature, precipitation, and evotranspiration influence river flows (Meuse and Rhine). When these flows fall below mandated environmental thresholds, surface water treatment plants must reduce operations or close, temporarily limiting available supply.
All four variables affect groundwater availability and recharge rates, potentially reducing the permitted extraction volumes of new sources in the future. For a more detailed description of surface water and groundwater source dynamics, see Section 2.1.3.
Solar radiation influences the long-term electricity generation potential of solar panels, as sustained higher irradiance levels lead to greater cumulative energy yields over time.
The review of the climate variables can be seen in Table LABEL:tbl:clim-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A. Note that instead of the evotranspiration variable, we directly provide the Standardized Precipitation-Evapotranspiration Index (SPEI) as it helps to quantify the severity, duration, and frequency of droughts by indicating whether conditions are wetter or drier than normal over a specific period.
Table 12: Climate model’s properties review.
Property Type Scope Unit
Temperature Dynamic Exogenous National °C
Precipitation Dynamic Exogenous National mm/day
Solar Radiation Dynamic Exogenous National W/m²
Stand. Precip.-Evap. Ind. (SPEI) Dynamic Exogenous National -
@@ 2.2.2 Energy System
The energy system in the BWF is modeled through three macro-drivers: electricity prices, greenhouse gas emission factor, and solar panels prices.
The electricity price is assumed to vary annually, influenced by inflation but subject to high uncertainty. On a higher resolution, prices also fluctuate hourly throughout the day. The electricity purchased by the grid has an associated greenhouse gas emissions factor, which can not be influcenced by the utilities but is adjusted yearly to account for the evolution of the national electricity generation mix.
As an alternative to grid electricity, participants may install solar photovoltaic panels. The unit cost of solar energy is modeled as a dynamic variable, changing annually to reflect improvements in the technology.
The capital investment associated with the installation of solar panels in water utility $w$ at year $y$ is:
[…]
where $v$ is the solar installation index within the set of photovoltaic systems in the water utility ($\mathcal{V}_{w}$), $\tau_{v}$ is the installation time, $\mathbf{1}_{\{\tau_{v}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), $c^{\text{solar}}(y)$ is the unit cost per kilowatt of capacity in year $y$, and $P_{v}$ is the nominal power capacity of the solar system $v$.
No cost is associated with the operation of the solar panels. These components have a fixed lifespan of 25 years, after which competitors must choose between investing in a replacement installation or reverting to full dependence on the electrical grid.
Table 13: Energy system model’s properties review.
Property Type Scope Unit
Electricity price Dynamic Exogenous National $\text{\texteuro}/kWh$
Emission factor Dynamic Exogenous National $tCO2eq/kWh$
Unit cost - solar panel Dynamic Endogenous National $\text{\texteuro}/kW$
@@ 2.2.3 Economy
2.2.3.1 Inflation Dynamics
Inflation captures the year-over-year change in the general price level. In the BWF, it affects all costs equally except energy prices, which follow different dynamics see 2.2.2.
As inflation compounds over time, it can substantially erode water utilities’ purchasing power and alter the financial viability of capital-intensive interventions. While inflation is an exogenous macroeconomic variable beyond the direct control of water utilities, competitors can adopt strategic responses: front-loading major investments to lock in current prices, or stress-testing their masterplan against multiple inflation scenarios to ensure robustness.
In the absence of reliable long-term inflation forecasts, a reasonable baseline assumption is the central bank’s target rate (e.g., 2% for the European Central Bank).
Table 14: Inflation’s properties review.
Property Type Scope Unit
Inflation rate Dynamic Exogenous National
Note that wherever unit prices are given only in base-year terms (e.g., only the value in year 2000 is reported in the input files), their value in year $y$ ($C(y)$) is obtained by multiplying by the cumulative inflation index
[…]
where $C(0)$ is the base-year unit price and $\pi(y)$ is the inflation rate in year $y$.
2.2.3.2 Budget Allocation
A national budget (predefined over the planning horizon) is distributed annually across utilities to fund infrastructure interventions and operational expenses.
Competitors must decide how to distribute this budget among utilities in every year. Budget distribution can follow several principled rules: proportional allocation to population size; inverse-proportional schemes favuoring smaller regions; income-based allocations reflecting socioeconomic considerations; inverse-proportional schemes favuoring less wealthy regions; or fully customised schemes.
Table 15: Budget allocation model’s properties review.
Property Type Scope Unit
National total budget Static National €/year
Budget allocation policy Option Water utility
2.2.3.3 Water Pricing
Water utilities generate revenue selling water to three types of customers: residential users, businesses, and, possibly, other water utilities. Residential and commercial customers are billed using a two-part pricing scheme: a fixed service charge ($\text{\texteuro}/year$) and a volumetric charge based on consumption ($\text{\texteuro}/m^{3}$). For simplicity, both retail customer types face identical rates within each utility (with also no differentiation by income class or other categories). Water transactions between utilities, however, use only a volumetric charge based on the net exchange at the end of the year.
Thus, the total revenue for a water utility $w$ in year $y$ is:
[…]
where $P_{w}^{\text{fixed}}(y)$ and $P_{w}^{\text{variable}}(y)$ are the fixed and volumetric components of the retail water price, $Q^{\text{BIL}}_{m}$ is the delivered billable demand in municipality $m$, $\mathcal{M}_{w}$ is the set of municipalities served by water utility $w$, $P_{w}^{\text{sell}}(y)$ is the volumetric charge that utility $w$ applies for inter-utility water sales, $Q^{w^{\prime}+}_{w}(y)$ is the net positive volume of water sold by utility $w$ to utility $w^{\prime}$, and $\mathcal{W}^{-}$ is the set of water utilities excluding utility $w$.
The net water exchange between utilities $w$ and $w^{\prime}$ is defined as:
$\Delta Q_{w}^{w^{\prime}}(y)$ $=\sum_{t\in\mathcal{Y}}\sum_{j\in\mathcal{J}_{(w,w^{\prime})}}Q_{j}(t)$ (17)
$Q_{w}^{w^{\prime}+}(y)$ $=\max(\Delta Q_{w}^{w^{\prime}}(y),0)$
$Q_{w}^{w^{\prime}-}(y)$ $=-\min(\Delta Q_{w}^{w^{\prime}}(y),0)$
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $Q_{j}(t)$ the flow over connection $j$ from the set of connections between the water utilities $\mathcal{J}_{(w,w^{\prime})}$ (with flow direction positive from $w$ to $w^{\prime}$).
Note that if a water utility has a negative net exchange with another utility (i.e., $\Delta Q_{w}^{w^{\prime}}(y)<0$), that will be regarded as a water import cost:
[…]
where $P_{w^{\prime}}^{\text{sell}}(y)$ is the volumetric charge applied by utility $w^{\prime}$ for water sales.
Participants must decide the water pricing adjustment strategy. They can either let all three quantities adjust according to inflation, or define a custom policy by specifying the yearly percentage increase for each of them independently.
Table 16: Water price model’s properties review.
Property Type Scope Unit
Fixed tariff component (service charge) Dynamic Endogenous Water utility €/year
Volumetric tariff component Dynamic Endogenous Water utility $\text{\texteuro}/m^{3}$
Water price for other utilities Dynamic Endogenous Water utility $\text{\texteuro}/m^{3}$
Water pricing adjustment policy Option Water utility
@@ 2.2.3.4 Bond Issuance
Whenever a water utility is unable to cover its expenditures in a given year, it finances the resulting deficit by issuing nationally backed bonds. Bonds are automatically generated to cover the utility debt in that year.
Specifically, the bond amount is $\kappa$ times the debt, i.e., $\mathrm{amount}_{i}=\kappa\cdot\mathrm{debt}_{w}(y)$ where $\kappa\in[1,2.5]$. A value of $\kappa$ equal to 1, implies that, depending on investor demand and prevailing market conditions, the proceeds from the bond issuance may barely cover the utility’s financing needs, leaving the balance for the following year around 0. In contrast, higher values of $\kappa$ (closer to 2.5) generate a cash surplus with the downside of a larger principal obligation to be repaid at maturity. Bonds are also characterised by a maturity of $M$ years, determining when the bond principal must be repaid, a coupon rate ($\mathrm{coupon}_{i}$), which determines the interest payments due each year, and a yield to maturity ($\mathrm{yield}_{i}$), which determines the price of the bond ($\mathrm{price}_{i}$).
Each year, the utility must repay the sum of principal amounts of all bonds reaching maturity plus the annual interest payments and and receives proceeds based on the bond price in that year. Formally:
$\text{PRI}_{w}(y)=\sum_{i\in\mathcal{B}_{w}(y):y=\tau_{i}+M}\mathrm{amount}_{i}$ (19)
$\text{INT}_{w}(y)=\sum_{i\in\mathcal{B}_{w}(y)}\mathrm{amount}_{i}\cdot\mathrm{coupon}_{i}$
[…]
where $i$ indicates the i-th bond, $\tau_{i}$ is the issuance year for bond $i$, $\mathcal{B}_{w}(y)$ is the set of bonds active for water utility $w$ in year $y$ and $\mathcal{B}_{w}(y):y=\tau_{i}+M$ the subset of bonds reaching maturity $M$.
The i-th bond’s coupon, yield, and price are given by:
$\mathrm{coupon}_{i}=r_{f}+\hat{\pi}(y=\tau_{i})$ (20)
$\mathrm{yield}_{i}=\mathrm{coupon}_{i}+a\cdot(1-d(y=\tau_{i}))$
$\mathrm{price}_{i}=\sum_{y=1}^{M}\frac{\mathrm{coupon}_{i}}{(1+\mathrm{yield}_{i})^{y}}+\frac{100}{(1+\mathrm{yield}_{i})^{M}}$
where $r_{f}$ is the risk-free rate (long-term government yield), $\hat{\pi}(y=\tau_{i})$ is the inflationary expectation at issuance year, $a$ is the sensitivity to investor demand, and $d(y=\tau_{i})$ is the uncertain demand factor for bond $i$ at issuance year.
Strong investor demand ($d(y)>1.0$) increases $\mathrm{PRO}_{w}(y)$ (through a lower yield and therefore higher price at issuance), while weak demand ($d(y)<1.0$) decreases it. This simulates real-world bond pricing where investor appetite introduces uncertainty to the utilities budgetting.
@@
While utilities cannot directly control bond yields, they can anticipate debt accumulation through scenario analysis and adopt strategies that maintain financial sustainability. For instance, utilities may design pacing of interventions to minimize borrowing, or evaluate alternative interventions that reduce the likelihood of large bond issuances during periods with unfavorable yield conditions.
The complete list of the bond model properties can be seen in Table LABEL:tbl:eb-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 17: Bonds model’s properties review.
Property Type Scope Unit
Balance Dynamic Endogenous Water utility
Bond amount to debt ratio Option National
Bond amount Static Bond
Bond issue date Static Bond
Bond maturity Static National years
Risk free rate Static National
Inflationary expectations Dynamic Exogenous National
Bond coupon Dynamic Endogenous Bond
Sensitivity Static National
Investors demand factor Dynamic Exogenous National
Bond yield Dynamic Endogenous Bond
Bond price Dynamic Endogenous Bond
2.3 System Levers
Participants make two types of strategic decisions: policies and interventions.